# I/O in Machine Learning Applications on HPC Systems: A 360-degree Survey

### v2

@@ 1. Introduction
Although many ML workloads still use the traditional I/O stack on HPC systems, some are shifting to use remote cloud computing systems that offer services such as data warehouses to store and retrieve large datasets. Cloud platforms such as Google Cloud Platform, AWS, and Microsoft Azure offer access to large datasets, including Google Analytics data (Google Cloud, 2024a) hosted on BigQuery (Google Cloud, 2024b), the $1000$ Genomes Project (Amazon Web Services, 2024a) hosted on Amazon Redshift (Amazon Web Services, 2024b), and NYC Taxi & Limousine Commission data (Microsoft Azure, 2024b) hosted on Azure Synapse Analytics (Microsoft Azure, 2024a). HPC workloads that aim to use these datasets require near-first-class integration with the cloud to enable efficient data transfers to HPC systems. When using the traditional I/O stack, high bandwidth and low latency data access are possible due to the locality of the data and the highly optimized stack. In contrast, access to data stored in a data warehouse is often accomplished through a network interface such as REST (Fielding and Taylor, 2000), SOAP (Box et al., 2000), or ODBC (Microsoft, 2024). Given the high latency inherent in accessing data stored in a remote storage location, optimizing data retrieval is crucial for reasonable performance. Despite the lower I/O performance imposed when using a data warehouse, they are still preferred in many scenarios because they provide a centralized data storage location, facilitate data integration from various sources, and enable data governance, which can result in higher-quality data. High-quality data is essential when training ML models because it directly influences their accuracy, reliability, and generalizability (et al, 2020c). In addition, the monetary cost needed to purchase and maintain an HPC system can far outweigh the cost of using managed cloud services (Carlyle et al., 2010; et al, 2013). Furthermore, integration of HPC systems with available cloud resources is necessary due to the cloud being used as a data storage solution from which datasets can be publicly shared and used by all.
@@
The topic of ML I/O and whether it has a significant role in poor performance has been a debate within the HPC community. Many point out that HPC systems often have enough node-local storage to cache the dataset, resulting in minimal time spent waiting on I/O (Schimmelpfennig et al., 2021; et al, 2021h, g, 2018d). Others mention that not all HPC systems have large amounts of node local storage or the datasets used are simply too large to be cached (Schimmelpfennig et al., 2021; et al, 2021h; Xu et al., 2021; et al, 2018d). In addition, due to the random sampling that commonly occurs during ML training, modern PFSs can often become an I/O bottleneck (Petersen, 2015; et al, 2020k, 2018d). Ultimately, whether I/O is a significant issue is subjective to the workload at hand. It depends on various factors such as dataset size, storage capacity, storage system bandwidth, I/O library implementation and their configurations, and I/O access patterns of applications. Furthermore, many ML application developers are not HPC storage experts; hence, even if their system is capable of providing great performance, transparent solutions such as caching and prefetching are needed to ensure the HPC systems are utilized effectively. Transparent solutions are of particular interest due to the increasing number of heterogeneous processors used within ML applications. These increase the complexity of optimally utilizing available compute resources. For these reasons, we believe that I/O and data management are key areas of interest to the ML community.

### v3

@@ 1. Introduction
Although many ML workloads still use the traditional I/O stack on HPC systems, some are shifting to use remote cloud computing systems that offer services such as data warehouses to store and retrieve large datasets. Cloud platforms such as Google Cloud Platform, AWS, and Microsoft Azure offer access to large datasets, including Google Analytics data (Google Cloud, 2024a) hosted on BigQuery (Google Cloud, 2024b), the $1000$ Genomes Project (Amazon Web Services, 2024a) hosted on Amazon Redshift (Amazon Web Services, 2024b), and NYC Taxi & Limousine Commission data (Microsoft Azure, 2024b) hosted on Azure Synapse Analytics (Microsoft Azure, 2024a). HPC workloads that aim to use these datasets require near-first-class integration with the cloud to enable efficient data transfers to HPC systems. When using the traditional I/O stack, high bandwidth and low latency data access are possible due to the locality of the data and the highly optimized stack. In contrast, access to data stored in a data warehouse is often accomplished through a network interface such as REST (Fielding, 2000), SOAP (Box et al., 2000), or ODBC (Microsoft, 2024). Given the high latency inherent in accessing data stored in a remote storage location, optimizing data retrieval is crucial for reasonable performance. Despite the lower I/O performance imposed when using a data warehouse, they are still preferred in many scenarios because they provide a centralized data storage location, facilitate data integration from various sources, and enable data governance, which can result in higher-quality data. High-quality data is essential when training ML models because it directly influences their accuracy, reliability, and generalizability (et al, 2020c). In addition, the monetary cost needed to purchase and maintain an HPC system can far outweigh the cost of using managed cloud services (Carlyle et al., 2010; et al, 2013). Furthermore, integration of HPC systems with available cloud resources is necessary due to the cloud being used as a data storage solution from which datasets can be publicly shared and used by all.
@@
The topic of ML I/O and whether it has a significant role in poor performance has been a debate within the HPC community. Many point out that HPC systems often have enough node-local storage to cache the dataset, resulting in minimal time spent waiting on I/O (Schimmelpfennig et al., 2021; Farrell et al., 2021; et al, 2021g, 2018c). Others mention that not all HPC systems have large amounts of node local storage or the datasets used are simply too large to be cached (Schimmelpfennig et al., 2021; Farrell et al., 2021; Xu et al., 2021; et al, 2018c). In addition, due to the random sampling that commonly occurs during ML training, modern PFSs can often become an I/O bottleneck (Petersen, 2015; et al, 2020i, 2018c). Ultimately, whether I/O is a significant issue is subjective to the workload at hand. It depends on various factors such as dataset size, storage capacity, storage system bandwidth, I/O library implementation and their configurations, and I/O access patterns of applications. Furthermore, many ML application developers are not HPC storage experts; hence, even if their system is capable of providing great performance, transparent solutions such as caching and prefetching are needed to ensure the HPC systems are utilized effectively. Transparent solutions are of particular interest due to the increasing number of heterogeneous processors used within ML applications. These increase the complexity of optimally utilizing available compute resources. For these reasons, we believe that I/O and data management are key areas of interest to the ML community.